# Deepfakes, Phrenology, Surveillance, and More! A Taxonomy of AI Privacy Risks

### v1

@@ 5.1. Theoretical extensions to the AI privacy risks taxonomy
There were three subcategories of privacy risk in Solove’s original taxonomy (Solove, 2006) for which we did not find relevant incidents in our dataset: Interrogation, Blackmail, and Decisional Interference.
@@
5.2. Mitigating privacy risks created and exacerbated by AI
Given that AI creates new types of privacy risks and exacerbates existing ones, it follows that there must be a need to provide practitioners with AI-specific product privacy guidance. Without this guidance, AI-created risks are unlikely to be addressed through standard operating procedures.
@@ 5.2. Mitigating privacy risks created and exacerbated by AI
Our taxonomy can serve as the first step to show how common AI requirements and capabilities map onto high-level privacy risks during the product development process. Future work can take an iterative approach, grounded on practitioners’ actual design needs, to model the translation function between AI product ideas and potential risks to consider. Indeed, envisioning with AI — i.e., treating AI as a product design material (Yang et al., 2018; Jansen and Colombo, 2023; Yildirim et al., 2022, 2023) — is an open and active area of research. Aligning with this line of research, future work can add to our taxonomy by systematizing AI capabilities and requirements, and the privacy risks they create and exacerbate, at a level of granularity that is useful for practitioners to ideate, communicate, and collaborate with product teams and stakeholders (Yildirim et al., 2022, 2023).
Additionally, future work can explore operationalizable approaches to understanding the utility-intrusiveness trade-offs when creating AI products (e.g., (Ernala et al., 2021)). For example, prior work in other AI-adjacent fields, such as Robotics, has explored how to correlate desired robot function with a minimally-invasive set of sensors (Eick and Anton, 2020). In a broader context of implementing privacy and security in software products, prior work has found that practitioners still largely see privacy and security in products as an “all or nothing” notion such that privacy comes at the expense of other important objectives (Tahaei et al., 2021; Gutfleisch et al., 2022). By providing practitioners with tools that help them find user-acceptable balances between the utility and intrusiveness of AI products and services, we hypothesize that practitioners will be able to better advocate and design for privacy (Tahaei et al., 2021).
Another promising avenue for future research is foregrounding tensions across data pipelines, practices, and stakeholders (i.e., data subjects, data observers, data beneficiaries, and data victims). By mirroring the first step in Rahwan’s Society in the Loop framework (Rahwan, 2018), AI practitioners can make concrete the envisioned value and the stakeholders of their proposed AI concepts. To assist in this process, future work can create artifacts that encourage practitioners to articulate the value proposition of their envisioned product, for whom that value is generated (i.e., data beneficiaries), whose data is processed to unlock that value (i.e., data subjects), who can be impacted by the data pipeline (i.e., data victims), and by which privacy risk (e.g., surveillance).
In practice, however, this type of early-stage discussion around AI utility and privacy risk can be challenging because: (i) practitioners do not necessarily understand the full potential and limitations of AI (Yildirim et al., 2022); (ii) privacy is often treated as compliance with general regulatory mandates rather than a product-specific design choice (Waldman, [n. d.]); and, (iii) practitioners do not have access to AI-specific tools that support their privacy works pertaining to the capabilities and requirements that AI brings to their products (Lee et al., 2024). Accordingly, there is a need for a greater understanding of where such tools and artifacts might be effectively incorporated into practitioners’ workflows.
5.3. Charting the design space for privacy-preserving AI/ML work
Our findings also broaden the design space for privacy-preserving AI and ML. For example, a recent meta-review of HAI principles and guidelines argues that privacy in ML-driven systems centers around the protection, control, and agency over personal data (Yurrita et al., 2022). Based on our findings, these criteria only consider some of the AI privacy risks we identified: they consider risks entailed by data requirements and processing, but do not consider the processing and dissemination risks newly created by AI. Similarly, while existing techniques for privacy-preserving AI/ML, such as Differential Privacy (Nasr et al., 2018) and Federated Learning (Li et al., 2020), offer quantifiable privacy guarantees, these guarantees only apply to some data processing risks and not the full range of risks we discuss in our taxonomy. For example, the “criminality classifier” that takes in photos of people’s faces and claims to predict their likelihood to be a criminal (Wu and Zhang, 2016) could be built with state-of-the-art differential privacy and through a federated learning architecture to “protect user data.” However, doing so will not address the physiognomy risk inherent to the idea itself, nor the exclusion and disclosure risks arising from how the data is collected and the inferences shared without consent.
@@ 5.4. Limitations
Our dataset is drawn from the AIAAIC database, and should not be interpreted as exhaustive and representative of all possible privacy risks created or exacerbated by AI technologies. Our goal in creating this taxonomy was to codify AI privacy risks based on an accounting of documented, real-world risks. To that end, AIAAIC is currently “the most comprehensive, detailed, and timely resource”131313https://www.aiaaic.org/aiaaic-repository/about-the-aiaaic-repository that is openly accessible and has been used by the community as the source to synthesize the harms caused by AI functionality (Raji et al., 2022). To mitigate the sampling bias introduced by our use of the AIAAIC, we tested the database’s coverage by independently collecting a list of 15 AI privacy incidents from various sources, e.g., social media posts, literature. Of the 15 incidents we collected, 13 were also included in AIAAIC. For the two incidents that were not included, we found very similar incidents in the database — i.e., similar privacy risks caused by the same technology (e.g., face recognition software) but of different products. As a comparison, we applied the same procedure to the AIID database and only found five incidents included. Thus, we believe that AIAAIC currently provides a pool of AI privacy accidents comprehensive enough for our goal.
We acknowledge that there will be a growing number of AI incidents, and that there may be existing AI incidents that were not captured in our dataset. For example, prior work has surfaced how algorithmic recommender systems can amplify embarrassing exposures through online social networks (Choi et al., 2015). Nevertheless, our taxonomy provides a solid foundation for understanding how the capabilities and requirements of AI change privacy risks. Since we ground our taxonomy on Solove’s taxonomy of privacy, which has remained highly influential and largely appropriate for nearly two decades, we are confident that our updated taxonomy can be flexibly adapted to encompass new risks as they arise.
@@ 6. Conclusion
In this paper, we conducted a systematic analysis of documented incidents of AI privacy risks to answer the question: How do modern advances in AI and ML change the privacy risks of a product or service? Our taxonomy, constructed from a corpus of 321 documented AI privacy incidents, reveals that while the incorporation of AI technologies into products does not necessarily change the privacy risks those products might entail, it often does. Our taxonomy reveals that AI can create new types of privacy risks when processing and disseminating end-user data. We showed, for example, that the unique capabilities of AI technologies (e.g., the ability to generate realistic but fake images) also create new types of privacy risks (e.g., exposure risks from deepfake pornography (Ayyub, 2018)). The taxonomy also reveals that the data requirements of AI technologies can exacerbate known privacy risks. For example, owing to the unique ability of AI to automatically identify individuals from low-fidelity images, governments are more motivated to capture facial images of all passengers that pass through major transportation hubs (e.g., (Fassler, 2021)). Our work suggests that AI-specific design guidance is needed for practitioners to negotiate the utility-intrusiveness trade-offs of AI-powered user experiences, and that many existing approaches to privacy-preserving machine learning (e.g., federated learning (Li et al., 2020)) address only a small subset of the unique privacy risks entailed by AI technologies.
@@ Acknowledgements.
This work was supported, in part, by the National Science Foundation through SaTC grants 2126058, 2126066, and 2316768.References
@@
- Harrison ([n. d.]) Maggie Harrison. [n. d.]. Startup Shocked When 4Chan Immediately Abuses Its Voice-Cloning AI. https://futurism.com/startup-4chan-voice-cloning-ai
@@
- Harwell (2019) Drew Harwell. 2019. A face-scanning algorithm increasingly decides whether you deserve the job. In Ethics of Data and Analytics. Auerbach Publications, 206–211.
@@
- Kelley et al. (2023) Patrick Gage Kelley, Celestina Cornejo, Lisa Hayes, Ellie Shuo Jin, Aaron Sedley, Kurt Thomas, Yongwei Yang, and Allison Woodruff. 2023. ”There will be less privacy, of course”: How and why people in 10 countries expect {AI} will affect privacy in the future. 579–603. https://www.usenix.org/conference/soups2023/presentation/kelley
- Kuo et al. (2023) Tzu-Sheng Kuo, Hong Shen, Jisoo Geum, Nev Jones, Jason I. Hong, Haiyi Zhu, and Kenneth Holstein. 2023. Understanding Frontline Workers’ and Unhoused Individuals’ Perspectives on AI Used in Homeless Services. https://doi.org/10.1145/3544548.3580882 arXiv:2303.09743 [cs].
- Langheinrich (2001) Marc Langheinrich. 2001. Privacy by Design — Principles of Privacy-Aware Ubiquitous Systems. In Ubicomp 2001: Ubiquitous Computing, Gerhard Goos, Juris Hartmanis, Jan van Leeuwen, Gregory D. Abowd, Barry Brumitt, and Steven Shafer (Eds.). Vol. 2201. Springer Berlin Heidelberg, Berlin, Heidelberg, 273–291. https://doi.org/10.1007/3-540-45427-6_23 Series Title: Lecture Notes in Computer Science.
@@
- Mikalef et al. (2022) Patrick Mikalef, Kieran Conboy, Jenny Eriksson Lundström, and Aleš Popovič. 2022. Thinking responsibly about responsible AI and ‘the dark side’ of AI. European Journal of Information Systems 31, 3 (May 2022), 257–268. https://doi.org/10.1080/0960085X.2022.2026621 Publisher: Taylor & Francis _eprint: https://doi.org/10.1080/0960085X.2022.2026621.
@@
- Nightingale and Farid (2022) Sophie J. Nightingale and Hany Farid. 2022. AI-synthesized faces are indistinguishable from real faces and more trustworthy. Proceedings of the National Academy of Sciences 119, 8 (Feb. 2022), e2120481119. https://doi.org/10.1073/pnas.2120481119 Publisher: Proceedings of the National Academy of Sciences.
@@
- Obar (2020) Jonathan A Obar. 2020. Sunlight alone is not a disinfectant: Consent and the futility of opening Big Data black boxes (without assistance). Big Data & Society 7, 1 (2020), 2053951720935615.
@@
- Simonite (2017) Tom Simonite. 2017. Facebook Can Now Find Your Face, Even When It’s Not Tagged. Wired (Dec. 2017). https://www.wired.com/story/facebook-will-find-your-face-even-when-its-not-tagged/ Section: tags.
@@
- Waldman ([n. d.]) Ari Ezra Waldman. [n. d.]. DESIGNING WITHOUT PRIVACY. HOUSTON LAW REVIEW ([n. d.]).
@@
- Wang and Kosinski (2018) Yilun Wang and Michal Kosinski. 2018. Deep neural networks are more accurate than humans at detecting sexual orientation from facial images. Journal of personality and social psychology 114, 2 (2018), 246.
@@
- Wickramasinghe et al. (2020) Chathurika S Wickramasinghe, Daniel L Marino, Javier Grandio, and Milos Manic. 2020. Trustworthy AI Development Guidelines for Human System Interaction. Proc. 13th Int. Conf. Hum. Syst. Interact (HSI), 130–136.

### v2

@@ 5.2. Theoretical extensions to the AI privacy risks taxonomy
We see our taxonomy as a living structure that helps scaffold the conversation about how advances in AI change privacy risks. But just as the capabilities and requirements of AI may change with future advances, so too might AI privacy risks. One way we might envision future AI privacy risks is by exploring the four subcategories of privacy risk in Solove’s original taxonomy (Solove, 2006) for which we did not find relevant incidents in our dataset: Interrogation, Blackmail, Breach of Confidentiality, and Decisional Interference. In the future, we may observe incidents where advances in AI meaningfully change or exacerbate these risks as well.
@@
Breach of Confidentiality
Breach of Confidentiality refers to an interpersonal risk between two people where one party discloses something to the other in confidence, and the other party violates this confidence by sharing it with third-parties (Solove, 2006). AI technologies could exacerbate the scale of this risk by enabling conversational agents capable of gaining users’ trust and guiding them to share sensitive information. For example, attackers can deploy such AI systems in high-stakes scenarios like healthcare and finance, and pose threats of breaching the confidentiality of the users by sharing the sensitive information they shared with the agent to third-parties.
@@
5.2.1. Future Work: Creating a living taxonomy of AI privacy risks
To our knowledge, our taxonomy is the first attempt to show how common AI requirements and capabilities map onto high-level privacy risks. As shown above, future AI privacy incidents can also expand the taxonomy. In addition, future AI privacy incidents may create new categories of privacy risk that go beyond Solove’s taxonomy (like the physiognomy risk we describe here). For example, many artists have been vocal about concerns about the theft of artistic style by generative AI (Knibbs, 2023). While these discussions currently center around notions of copyright and intellectual property, we can envision new types of privacy risk as well: e.g., artistic styles might contain personally identifiable or sensitive information. We envision that our taxonomy can complement ongoing crowd-sourced efforts at curating and organizing AI incidents such as the AIAAIC (Pownall, 2023) and AIID181818https://incidentdatabase.ai/ by providing a framework to formally synthesize and identify emerging privacy risks in AI incidents. With that in mind, the research team is building a website191919The website will be available at https://privacytaxonomy.ai/ and https://aiprivacytaxonomy.com/ to present our taxonomy of AI privacy risks, and is also planning to expand this website to collect and aggregate submissions of new incidents related to these risks.
@@ 5.2.1. Future Work: Creating a living taxonomy of AI privacy risks
To present the AI privacy taxonomy in forms useful to the HCI and AI communities, future work can take an iterative approach, grounded on practitioners’ and academics’ actual design and research needs, to model the translation function between AI technology ideas and potential risks to consider. Indeed, envisioning with AI — i.e., treating AI as a design material (Yang et al., 2018; Jansen and Colombo, 2023; Yildirim et al., 2022, 2023) — is an open and active area of research. Aligning with this line of research, future work can add to our taxonomy by systematizing AI capabilities and requirements, and the privacy risks they create and exacerbate, at a level of granularity that is useful for practitioners and researchers to ideate, communicate, and collaborate with product teams and stakeholders (Yildirim et al., 2022, 2023).
@@ 5.3. Limitations
We consciously took an “incident-based” approach when constructing our taxonomy. There is a great deal of hype about what AI technologies can do, blurring the lines between speculation and reality (Kapoor and Narayanan, Arvind, 2023). The overabundance of speculative risks necessitated that we limit our consideration to those that journalists and the public-at-large have recognized as harmful as chronicled in the AIAAIC database. With that in mind, our dataset should not be interpreted as inclusive and representative of every possible privacy risk created or exacerbated by AI technologies: it is a repository of many privacy risks that have been realized in practice.
Our goal in creating this taxonomy was to codify AI privacy risks based on an accounting of documented, real-world risks. To that end, AIAAIC is currently “the most comprehensive, detailed, and timely resource”202020https://www.aiaaic.org/aiaaic-repository/about-the-aiaaic-repository that is openly accessible and has been used by the community as the source to synthesize the harms caused by AI functionality (Raji et al., 2022). To mitigate the sampling bias introduced by our use of the AIAAIC, we tested the database’s coverage by independently collecting a list of 15 AI privacy incidents from various sources, e.g., social media posts, literature. Of the 15 incidents we collected, 13 were also included in AIAAIC. For the two incidents that were not included, we found very similar incidents in the database — i.e., similar privacy risks caused by the same technology (e.g., face recognition software) but of different products. As a comparison, we applied the same procedure to the AIID database and only found five incidents included. Thus, we believe that AIAAIC currently provides a pool of AI privacy accidents comprehensive enough for our goal.
We acknowledge that there will be a growing number of AI incidents, and that there may be existing AI incidents that were not captured in our dataset. For example, prior work has surfaced how algorithmic recommender systems can amplify embarrassing exposures through online social networks (Choi et al., 2015). Nevertheless, our taxonomy provides a solid foundation for understanding how the capabilities and requirements of AI change privacy risks. Since we ground our taxonomy on Solove’s taxonomy of privacy, which has remained highly influential and largely appropriate for nearly two decades, we are confident that our updated taxonomy can be flexibly adapted to encompass new risks if and as they are realized beyond academic inquiry.
Finally, we acknowledge that “privacy” is a broad and context-dependent concept that is susceptible to biased interpretation based on the research team’s background. We are an interdisciplinary research team with diverse expertise across HCI, AI, security and privacy, policy, and design. We mitigated the potential for bias by: (i) building our taxonomy on top of Solove’s existing and widely accepted taxonomy; (ii) ensuring that multiple coders independently agreed on the risks entailed (or not) by a specific incident; and, (iii) dutifully analyzing all incidents, in the AIAAIC database, that were independently characterized by people outside of our research as being privacy-pertinent.
@@ 6. Conclusion
In this paper, we conducted a systematic analysis of documented incidents of AI privacy risks to answer the question: How do modern advances in AI and ML change privacy risks? Our taxonomy, constructed from a corpus of 321 documented AI privacy incidents, reveals that while the incorporation of AI technologies into products does not necessarily change the privacy risks those products might entail, it often does. Our taxonomy reveals that AI can create new types of privacy risks when processing and disseminating end-user data. We showed, for example, that the unique capabilities of AI technologies (e.g., the ability to generate realistic but fake images) also create new types of privacy risks (e.g., exposure risks from deepfake pornography (Ayyub, 2018)). The taxonomy also reveals that the data requirements of AI technologies can exacerbate known privacy risks. For example, owing to the unique ability of AI to automatically identify individuals from low-fidelity images, governments are more motivated to capture facial images of all passengers that pass through major transportation hubs (e.g., (Fassler, 2021)). Our work suggests that AI-specific design guidance is needed for practitioners to negotiate the utility-intrusiveness trade-offs of AI-powered user experiences, and that many existing approaches to privacy-preserving machine learning (e.g., federated learning (Li et al., 2020)) address only a small subset of the unique privacy risks entailed by AI technologies.
@@ Acknowledgements.
This work was generously funded, in part, by NSF SaTC grants #2316768 and #2126066. We thank our anonymous reviewers for their helpful comments in revising this paper. We also thank the contributors and maintainers of the AIAAIC incident database.References
@@
- Acquisti et al. (2015) Alessandro Acquisti, Laura Brandimarte, and George Loewenstein. 2015. Privacy and human behavior in the age of information. Science 347, 6221 (2015), 509–514.
@@
- Albert and Delano (2022) Kendra Albert and Maggie Delano. 2022. Algorithmic Exclusion. https://doi.org/10.2139/ssrn.4122529
@@
- Andrus and Villeneuve (2022) McKane Andrus and Sarah Villeneuve. 2022. Demographic-Reliant Algorithmic Fairness: Characterizing the Risks of Demographic Data Collection in the Pursuit of Fairness. 2022 ACM Conference on Fairness, Accountability, and Transparency (2022).
@@
- Cyphers (2021) Bennett Cyphers. 2021. Google’s FLoC Is a Terrible Idea. https://www.eff.org/deeplinks/2021/03/googles-floc-terrible-idea
- Das et al. (2018) Sauvik Das, Joanne Lo, Laura Dabbish, and Jason I. Hong. 2018. Breaking! A Typology of Security and Privacy News and How It’s Shared. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems. ACM, Montreal QC Canada, 1–12. https://doi.org/10.1145/3173574.3173575
@@
- doteveryone (2020) doteveryone. 2020. Consequence Scanning – an agile practice for responsible innovators – doteveryone. https://doteveryone.org.uk/project/consequence-scanning/
@@
- Harrison (2023) Maggie Harrison. 2023. Startup Shocked When 4Chan Immediately Abuses Its Voice-Cloning AI. https://futurism.com/startup-4chan-voice-cloning-ai
@@
- Holohan et al. (2019) Naoise Holohan, Stefano Braghin, Pól Mac Aonghusa, and Killian Levacher. 2019. Diffprivlib: the IBM differential privacy library. ArXiv e-prints 1907.02444 [cs.CR] (July 2019).
@@
- Kelley et al. (2023) Patrick Gage Kelley, Celestina Cornejo, Lisa Hayes, Ellie Shuo Jin, Aaron Sedley, Kurt Thomas, Yongwei Yang, and Allison Woodruff. 2023. ”There will be less privacy, of course”: How and why people in 10 countries expect {AI} will affect privacy in the future. In Nineteenth Symposium on Usable Privacy and Security (SOUPS 2023). 579–603. https://www.usenix.org/conference/soups2023/presentation/kelley
- Knibbs (2023) Kate Knibbs. 2023. A New Tool Helps Artists Thwart AI—With a Middle Finger. Wired (Oct. 2023). https://www.wired.com/story/kudurru-ai-scraping-block-poisoning-spawning/
- Kokolakis (2017) Spyros Kokolakis. 2017. Privacy attitudes and privacy behaviour: A review of current research on the privacy paradox phenomenon. Computers & Security 64 (2017), 122–134. https://doi.org/10.1016/j.cose.2015.07.002
- Kuo et al. (2023) Tzu-Sheng Kuo, Hong Shen, Jisoo Geum, Nev Jones, Jason I. Hong, Haiyi Zhu, and Kenneth Holstein. 2023. Understanding Frontline Workers’ and Unhoused Individuals’ Perspectives on AI Used in Homeless Services. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. ACM, Hamburg Germany, 1–17. https://doi.org/10.1145/3544548.3580882
- Langheinrich (2001) Marc Langheinrich. 2001. Privacy by Design — Principles of Privacy-Aware Ubiquitous Systems. In Ubicomp 2001: Ubiquitous Computing, Gerhard Goos, Juris Hartmanis, Jan van Leeuwen, Gregory D. Abowd, Barry Brumitt, and Steven Shafer (Eds.). Vol. 2201. Springer Berlin Heidelberg, Berlin, Heidelberg, 273–291. https://doi.org/10.1007/3-540-45427-6_23
@@
- Meurisch and Mühlhäuser (2022) Christian Meurisch and Max Mühlhäuser. 2022. Data Protection in AI Services: A Survey. Comput. Surveys 54, 2 (March 2022), 1–38. https://doi.org/10.1145/3440754
@@
- Mikalef et al. (2022) Patrick Mikalef, Kieran Conboy, Jenny Eriksson Lundström, and Aleš Popovič. 2022. Thinking responsibly about responsible AI and ‘the dark side’ of AI. European Journal of Information Systems 31, 3 (May 2022), 257–268. https://doi.org/10.1080/0960085X.2022.2026621
@@
- Murakonda and Shokri (2020) Sasi Kumar Murakonda and Reza Shokri. 2020. ML Privacy Meter: Aiding regulatory compliance by quantifying the privacy risks of machine learning. arXiv preprint arXiv:2007.09339 (2020).
@@
- Nightingale and Farid (2022) Sophie J. Nightingale and Hany Farid. 2022. AI-synthesized faces are indistinguishable from real faces and more trustworthy. Proceedings of the National Academy of Sciences 119, 8 (Feb. 2022), e2120481119. https://doi.org/10.1073/pnas.2120481119
@@
- Obar (2020) Jonathan A Obar. 2020. Sunlight alone is not a disinfectant: Consent and the futility of opening Big Data black boxes (without assistance). Big Data & Society 7, 1 (2020), 2053951720935615. https://doi.org/10.1177/2053951720935615
@@
- Simonite (2017) Tom Simonite. 2017. Facebook Can Now Find Your Face, Even When It’s Not Tagged. Wired (Dec. 2017). https://www.wired.com/story/facebook-will-find-your-face-even-when-its-not-tagged/
@@
- Stark and Hutson (2021) Luke Stark and Jevan Hutson. 2021. Physiognomic artificial intelligence. Fordham Intell. Prop. Media & Ent. LJ 32 (2021), 922.
@@
- Waldman (2017) Ari Ezra Waldman. 2017. Designing Without Privacy. Houston Law Review (March 2017). https://papers.ssrn.com/abstract=2944185
@@
- Wang and Kosinski (2018) Yilun Wang and Michal Kosinski. 2018. Deep neural networks are more accurate than humans at detecting sexual orientation from facial images. Journal of personality and social psychology 114, 2 (2018), 246. https://doi.org/10.1037/pspa0000098
@@
- Wickramasinghe et al. (2020) Chathurika S. Wickramasinghe, Daniel L. Marino, Javier Grandio, and Milos Manic. 2020. Trustworthy AI Development Guidelines for Human System Interaction. In 2020 13th International Conference on Human System Interaction (HSI). IEEE, Tokyo, Japan, 130–136. https://doi.org/10.1109/HSI49210.2020.9142644
@@
- Windl et al. (2023) Maximiliane Windl, Verena Winterhalter, Albrecht Schmidt, and Sven Mayer. 2023. Understanding and Mitigating Technology-Facilitated Privacy Violations in the Physical World. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (Hamburg, Germany). Association for Computing Machinery, New York, NY, USA, Article 585, 16 pages. https://doi.org/10.1145/3544548.3580909
- Wong et al. (2023) Richmond Y. Wong, Michael A. Madaio, and Nick Merrill. 2023. Seeing Like a Toolkit: How Toolkits Envision the Work of AI Ethics. Proceedings of the ACM on Human-Computer Interaction 7, CSCW1 (April 2023), 1–27. https://doi.org/10.1145/3579621
@@
Appendix A Appendix
| Privacy Risk | Cohen’s Kappa |
| Surveillance | 0.88 |
| Identification | 0.94 |
| Secondary Use | 0.84 |
| Aggregation | 0.87 |
| Phrenology / Physiognomy | 1 |
| Exclusion | 0.90 |
| Insecurity | 0.94 |
| Exposure | 1 |
| Disclosure | 1 |
| Distortion | 1 |
| Increased Accessibility | 1 |
| Intrusion | 0.94 |
| Average | 0.94 |